# Learning Japanese with Jouzu: Interaction Outcomes with Stylized Dialogue Fictional Agents

### v1

@@ Abstract.
This study investigates how stylized, voiced agents shape user interaction in a multimodal language learning environment. We conducted a mixed-methods evaluation of 54 participants interacting with anime-inspired characters powered by large language models and expressive text-to-speech synthesis. These agents responded in Japanese character language, offering users asynchronous, semi-structured conversation in varying speech styles and emotional tones. We analyzed user engagement patterns, perceived usability, emotional responses, and learning behaviors, with particular attention to how agent stylization influenced interaction across language proficiency levels and cultural backgrounds. Our findings reveal that agent design, especially voice, persona, and linguistic style, substantially affected user experience, motivation, and strategy. This work contributes to the understanding of affective, culturally stylized agents in human–agent interaction and offers guidance for designing more engaging, socially responsive systems.
@@
copyright: noneconference: ; ;
Figure 1. A screenshot from our Japanese language learning app, Jouzu, featuring “Hikari,” an anime-inspired conversational agent. Users interact via text (”send”) or voice (”record”), receiving stylized, in-character responses designed to enhance engagement and cultural immersion.
[…]
Screenshot of a conversation between the user and an anime-style agent named Hikari. The user asks for a Japanese word, and the agent replies in Japanese with a stylized, character-specific response.
@@ 1. Introduction
As conversational agents become integrated into everyday tools, the stylization of an agent’s voice, character, and aesthetics have become essential to the user experience. In human–agent interaction (HAI), conversational agents are not just tools, but social actors. They express personality, mediate emotion, and structure interaction in ways that influence user trust, engagement, and learning. This paper examines how stylistic and affective features of Japanese learning agents, particularly conversational characters, impact user interaction with the agents and user’s Japanese learning outcomes.
@@
We focus on a multimodal interface where users engage with fictional agents designed to emulate recognizable character archetypes from anime, which is a form of Japanese animation that has recently become ingrained in pop culture around the world. By leveraging large language models (LLMs) and an expressive multilingual text-to-speech (TTS) system using Style-BERT-VITS2 (Rackauckas and Hirschberg, 2025), we produce emotionally rich, character-specific Japanese text and speech responses in conversational character agents. Unlike neutral chatbots, these agents employ language with non-standard pronouns, stylistic particles, and expressive phrasing to simulate social presence and emotional realism (Kinsui and Yamakido, 2015). Through these features, the agents act as affective, narrative participants who provide an amusing and engaging conversational environment rather than a traditional, robotic chatbot tutor.
Jouzu [^1] https://jouzu.ai features not only conversational characters, but also prescripted question-answer (QA) dialogues, or lessons, an in-text sentence-level inspector that defines and provides the readings of Japanese words, and flashcards to practice words learned in the lessons and conversational environments.
@@
In a mixed-methods study with 54 participants, we evaluate how users of varying proficiency and linguistic backgrounds interacted with this application. We collected quantitative ratings of usability, engagement, and agent perception, alongside open-ended feedback and behavioral data from chatbot logs. In this study, we analyze how agent stylization such as voice, visual design, and character behavior modulates user interaction, emotional response, and task engagement. This contributes to ongoing discourse in HAI on the role of personality, culture, and embodiment in agent design.
2. Referenced Works
@@
Language learning systems increasingly incorporate conversational agents to simulate spoken practice in a low-risk environment. Traditional computer-assisted language learning (CALL) tools have relied on scripted responses and task-based exercises, but newer systems use LLMs to support open-ended dialogue, generate feedback, and scaffold instruction (Cao and Zhong, 2023; Dornburg and Davin, 2024). Question-answer chatbots have been deployed in multiple languages to provide real-time practice, although most remain text-based (Li et al., 2022). More recent work has incorporated empathetic language to reduce anxiety and increase user retention (Siyan et al., 2024a, b). Aiba et al. (Aiba et al., 2024) introduced a system in which English learners could practice conference-style questions with GPT-generated speech, suggesting that combining TTS with LLMs encourages natural spoken interaction.
Voiced characters extend this trend by merging LLM dialogue with expressive speech synthesis and cultural stylization. While most prior systems employ neutral or utilitarian speech, Jouzu introduces character-specific language rooted in anime tropes—offering linguistic diversity and affective realism. These characters use “character language” featuring stylized pronouns, emphatic particles, and emotionally expressive phrasing, contributing to both social presence and learner motivation.
@@
Kinsui and Yamakido (Kinsui and Yamakido, 2015) describe “character language” as a stylized form of Japanese used in fiction to signal personality, gender, or origin. These features include non-standard pronouns, archaic verb forms, and distinctive particles. Although rare in textbooks, they are recognizable and emotionally charged for many learners. Li et al. (Li et al., 2023) explored this concept with ChatHaruhi, a chatbot modeled on anime characters, but did not include voice interaction.
In contrast, Jouzu introduces voiced agents with character-specific linguistic stylization rooted in anime tropes. These agents use “character language” (Kinsui and Yamakido, 2015) to emulate fictional personas with affective realism. This represents a shift from purely pedagogical interactions toward immersive, socially engaging dialogue that blends linguistic and cultural learning.
@@
Multimodal interfaces are central to current human-computer interaction (HCI) discussions about language learning. OpenAI’s GPT-4o (et al., 2024), for example, enables faster response generation with integrated audio (Ying et al., 2024), supporting smooth, responsive dialogue. At the same time, systems must address concerns about equity, safety, and engagement. Chan and Tsi (Chan and Tsi, 2023) note that computer-assisted tutors must maintain educational integrity while sustaining user interest. Xue et al. (Xue et al., 2023) highlight the importance of filtering harmful outputs in generative systems.
@@
Previous HCI research has leveraged user studies to investigate and proposed guidelines for interactions with learner dialogue systems. For example, through a user study of 107 users, Koyoturk et al. assert the need for effective prompting in learner chatbot interactions, highlighting the need to provide user guidelines to learn how to prompt the agents (Koyuturk et al., 2025). Cai et al. conducted a user study with a group of three undergraduate students to advance dialogue system understanding in small group learning (Cai et al., 2024). Annamalai et al. conducted a user study with 360 undergraduate students to investigate chatbot language learning in Malaysian higher education institutions (Annamalai et al., 2023). Finally, Li et al. conducted a user study with 24 students to assess a curriculum-driven chatbot to match students’ classroom curricula (Li et al., 2024). Similarly, we employ a user study with 52 users to analyze user interactions and learning outcomes after their use of Jouzu, centered around the conversational character agents.
@@ 3. Method
A user study was conducted with 54 participants initially recruited; of these, 52 fully completed all parts of the study, including surveys and app usage. The analysis below focuses on these 52 complete cases to ensure consistency and validity. We examine app usability, learning outcomes, engagement, impact of stylized characters on the learning experience, and to compare Jouzu to other language learning apps like Duolingo. In addition, we gathered qualitative feedback on user experiences and analyzed interaction patterns between users and conversational agents.
Participants were recruited from a university computer science class. They were then instructed to freely use the app for at least one hour over a period of two weeks. In addition, we asked them to use the conversational agents present in the app. Users completed a survey consisting of multiple choice questions and open responses before (see Appendix A) and after the app (see Appendix B). The initial survey collected demographic and background information from users and presented an optional 10-question multiple choice Japanese quiz with an optional short paragraph response in Japanese (see Appendix C). The post-use survey asked several questions about the user experience, followed by a new 10-question Japanese quiz of the same difficulty level and an open response in Japanese (see Appendix C). Participants were rewarded with 5 assignment points in their university class. Of the 54 participants, 52 fully completed the task.
@@ 3.2. Hypotheses and Analysis Design
To examine potential cross-linguistic differences in user experience, participants were grouped by native language: Chinese and non-Chinese speakers.
To assess the significance of these cross-linguistic differences, we conducted independent-sample t-tests comparing mean group ratings between Chinese and non-Chinese participants across all six experience categories.
We also grouped users by self-reported Japanese proficiency (Complete Beginner, Beginner, Intermediate, Native) to analyze mean ratings by language experience level. A heatmap was generated to visualize rating trends across these categories.
@@ 3.4.1. Multiple Choice
Users were asked ten multiple choice questions before and after use of the app. All questions were optional. The questions are included in . Pre- and post-use completion rates and accuracy scores were compared to assess realized learning outcomes in terms of grammar and vocabulary.
Ten questions were asked. The first five questions included vocabulary sampled from the characters’ individual language styles. The next five were questions sampled from each Japanese Language Proficiency Test Level (N1 through N5).
The pre- and post-tests were different, but were created to be the same difficulty for respondents.
3.4.2. Open Paragraph
Participants were invited to write a short paragraph in Japanese both before and after using the app. This task was optional and open-ended, allowing users at any level to attempt expressive written output. Pre- and post-use submissions were compared for fluency, vocabulary diversity, and structural complexity.

### v2

# Animating Language Practice: Engagement with Stylized Conversational Agents in Japanese Learning

@@ Abstract.
We explore Jouzu, a Japanese language learning application that integrates large language models with anime-inspired conversational agents. Designed to address challenges learners face in practicing natural and expressive dialogue, Jouzu combines stylized character personas with expressive text-to-speech to create engaging conversational scenarios. We conducted a two-week in-the-wild deployment with 52 Japanese learners to examine how such stylized agents influence engagement and learner experience. Our findings show that participants interacted frequently and creatively, with advanced learners demonstrating greater use of expressive forms. Participants reported that the anime-inspired style made practice more enjoyable and encouraged experimenting with different registers. We discuss how stylization shapes willingness to engage, the role of affect in sustaining practice, and design opportunities for culturally grounded conversational AI in computer-assisted language learning (CALL). By framing our findings as an exploration of design and engagement, we highlight opportunities for generalization beyond Japanese contexts and contribute to international HCI scholarship.
@@
Figure 1. A screenshot from our Japanese language learning mobile app, Jouzu, featuring “Hikari,” an anime-inspired conversational agent. Users interact via text (”send”) or voice (”record”), receiving stylized, in-character responses designed to enhance engagement and cultural immersion.:
@@ 1. Introduction
As conversational agents become integrated into everyday tools, the stylization of an agent’s voice, character, and aesthetics has become essential to the user experience. In human–computer interaction (HCI), conversational agents are not just tools, but social actors. They express personality, mediate emotion, and structure interaction in ways that influence user trust, engagement, and learning. Stylization is therefore a central design dimension with implications not only for education, but also for how people engage with technology in culturally specific and affectively rich ways across diverse domains.
At the same time, learning Japanese presents challenges reflected in both recent and foundational scientific literature. For example, it is ranked among the most difficult languages for native English speakers to learn, requiring approximately 2,200 classroom hours to reach proficiency (usdos_fsi_foreign_language_training). Beyond its multiple writing systems, Japanese diverges significantly from English in grammar (boboc2023) and phonology (ohata2004). Traditional tools such as textbooks are often rigid, impersonal, and lack spoken interaction, contributing to learner disengagement and high dropout rates. Conversational practice remains one of the most effective techniques for language acquisition, yet it is often limited by social anxiety and access to native speakers (li2022). These constraints position language learning as a high-barrier context in which the role of agent stylization may be particularly impactful.
@@
We focus on providing a multimodal interface where Japanese learners engage with fictional agents designed to emulate recognizable character archetypes from anime, a form of Japanese animation that has become ingrained in global pop culture. To provide fun, engaging, and personalized learning to students who may otherwise quit due to the high barriers to entry of Japanese, we conducted an in-the-wild study using the Jouzu mobile app, which leverages large language models (LLMs) and an expressive multilingual text-to-speech (TTS) system using Style-BERT-VITS2 (rackauckas2025benchmarkingexpressivejapanesecharacter). Unlike prior computer-assisted language learning (CALL) systems that rely on neutral chatbots or scripted tutors, produces emotionally rich, character-specific Japanese text and speech responses in conversational agents. These agents employ language with non-standard pronouns, stylistic particles, and expressive phrasing to simulate social presence and emotional realism (kinsui2015). They are paired with anime-inspired visuals and animations to create an environment that combines cultural familiarity with affective depth. Through these features, the agents act as narrative participants who provide a stress-free and engaging conversational environment rather than a traditional, robotic chatbot tutor.
As a whole, Jouzu combines such conversational characters with prescripted QA dialogues, referred to as ”lessons,” an in-text sentence-level inspector that provides word-by-word definitions and readings and flashcards to practice vocabulary from both lessons and conversation created with the in-text inspector. These multimodal conversational environments are designed to replicate human interaction without judgment, thereby encouraging playful exploration in chats alongside structured practice in lessons and flashcards. The visuals of the app, including flashcards, are grounded in anime stylization, tying the system’s linguistic playfulness to a coherent aesthetic world.
@@
In a mixed-methods study with 52 participants, we investigated how users of varying proficiency and linguistic backgrounds interacted with the application. We collected quantitative ratings of usability, engagement, and agent perception, alongside open-ended feedback and behavioral data from chatbot logs. Our analysis describes patterns in how agent stylization including voice, visual design, and character behavior related to reported user experience, emotional responses, and interaction strategies. In some cases, we also tested for significant differences between user subgroups (e.g., native language and proficiency level), allowing us to examine potential impacts of linguistic background on app usage. This study therefore not only contributes to the design of language learning systems, but also extends ongoing discourse in HAI and HCI on the role of personality, culture, and embodiment in agent design.
We have three contributions:
  1. (1)
    A novel conversational learning system that combines large language models, a style-transfer TTS engine (Style-BERT-VITS2), and anime-inspired personas to support second-language (L2) practice.
  2. (2)
    An in-the-wild deployment study with 52 Japanese language learners, offering empirical insights into how stylized conversational agents influence learner engagement, motivation, and willingness to experiment with language in everyday contexts.
  3. (3)
    Design implications for HCI and CALL, highlighting how cultural stylization and affective expression can support engagement and creative language use, and discussing the opportunities and challenges of deploying such systems across different learner backgrounds and proficiency levels.
What distinguishes this work is its exploration of stylization as more than surface aesthetics, examining how anime-inspired characters, expressive speech synthesis, and culturally situated design elements interact with learners across diverse backgrounds. More broadly, this study contributes to HCI by showing how culturally resonant, affective stylization can be leveraged to design agents that are not only functional but also socially and emotionally meaningful. This positions stylization as a key lever for designing next-generation conversational agents in HCI, where personality and culture are as central as functionality.
2. Related Work
Early HCI research established principles that remain directly relevant to conversational agents for language learning. Card, Moran, and Newell’s keystroke level model provided a rigorous framework for predicting expert, error free task performance and showed how even small interaction costs shape usability (Card1980). While language learners are novices rather than experts, this insight underscores the need to minimize friction in conversational systems. Furnas et al. highlighted the vocabulary problem and demonstrated that users often employ diverse and mismatched terms, a challenge that remains central to natural language interfaces and learner input variability (Furnas1987). Grudin emphasized that computing is inherently social and collaborative and framed systems as situated within organizational and cultural contexts. This insight resonates with the social nature of language learning (Grudin1994). Shneiderman articulated core usability principles of consistency, informative feedback, and error prevention and handling in his book Designing the User Interface. These principles continue to underpin effective CALL design (Shneiderman1986). Our work extends these foundations by exploring how stylized voiced characters draw on cultural and affective dimensions to support engagement and learning in multimodal CALL environments.
@@ 2. Related Work
Language learning systems increasingly incorporate conversational agents to simulate spoken practice in a low-risk environment. Traditional computer-assisted language learning (CALL) tools have relied on scripted responses and task-based exercises, but newer systems use LLMs to support open-ended dialogue, generate feedback, and scaffold instruction (cao2023; dornburg2024). Question-answer chatbots have been deployed in multiple languages to provide real-time practice, although most remain text-based (li2022). More recent work has incorporated empathetic language to reduce anxiety and increase user retention (siyan2024a; siyan2024b). Aiba et al. (aiba2024) introduced a system in which English learners could practice conference-style questions with GPT-generated speech, suggesting that combining TTS with LLMs encourages natural spoken interaction. Voiced characters extend this trend by merging LLM dialogue with expressive speech synthesis and cultural stylization. While most prior systems employ neutral or utilitarian speech, Jouzu introduces character-specific language rooted in anime tropes –offering linguistic diversity and affective realism. These characters use “character language” featuring stylized pronouns, emphatic particles, and emotionally expressive phrasing, contributing to both social presence and learner motivation.
@@
Kinsui and Yamakido (kinsui2015) describe “character language” as a stylized form of Japanese used in fiction to signal personality, gender, or origin. These features include non-standard pronouns, archaic verb forms, and distinctive particles. Although rare in textbooks, they are recognizable and emotionally charged for many learners. Li et al. (li2023haruhi) explored this concept with ChatHaruhi, a chatbot modeled on anime characters, but did not include voice interaction. In contrast, Jouzu introduced voiced agents with character-specific linguistic stylization rooted in anime tropes. These agents use “character language” (kinsui2015) to emulate fictional personas with affective realism. This represents a shift from purely pedagogical interactions toward immersive, socially engaging dialogue that blends linguistic and cultural learning.
@@
Multimodal interfaces are central to current human-computer interaction (HCI) discussions about language learning. OpenAI’s GPT-4o (openai2024gpt4technicalreport), for example, enables faster response generation with integrated audio (ying2024), supporting smooth, responsive dialogue. At the same time, systems must address concerns about equity, safety, and engagement. Chan and Tsi (chan2023) note that computer-assisted tutors must maintain educational integrity while sustaining user interest. Xue et al. (xue2023) highlight the importance of filtering harmful outputs in generative systems.
@@
Previous HCI research has leveraged user studies to investigate and proposed guidelines for interactions with learner dialogue systems. For example, through a user study of 107 users, Koyoturk et al. assert the need for effective prompting in learner chatbot interactions, highlighting the need to provide user guidelines to learn how to prompt the agents (koyuturk2025understandinglearnerllmchatbotinteractions). Cai et al. conducted a user study with a group of three undergraduate students to advance dialogue system understanding in small group learning (caietal). Annamalai et al. conducted a user study with 360 undergraduate students to investigate chatbot language learning in Malaysian higher education institutions (Annamalai2023Chatbots). Finally, Li et al. conducted a user study with 24 students to assess a curriculum-driven chatbot to match students’ classroom curricula (li-etal-2024-curriculum). Similarly, we employ a user study with 52 users to analyze user interactions and learning outcomes after their use of Jouzu, centered around the conversational character agents.
3. System Architecture
Jouzu integrates voiced character agents, scripted dialogues, low-stakes stylized quizzes, and an interactive sentence-level inspector for building vocabulary flashcards. While flashcards and scripted lessons are familiar features in computer-assisted language learning, Jouzu’s design emphasizes stylization and immersion, elevating these methods into a playful and affectively engaging environment.
3.1. Original Characters
Jouzu features three original characters with distinct personalities, backstories, voices, and speaking styles. Each character was designed to represent recognizable anime tropes. In this study, the three characters deployed were a young girl attending a magic school who aspires to master light magic (Hikari), a Japanese kitsune deity capable of taking human form (Kitsune), and a samurai from Edo-period Japan. Each character employs a distinctive character language style (i.e., a casual young female, a wise centuries-old female deity, and a masculine-coded samurai), providing learners with exposure to diverse forms of Japanese vocabulary, grammar, and expressive speech.
These characters coexist within a stylized fictional world that blends elements of magic, prayer, and samurai culture. Professional character and background art reinforce immersion, grounding the linguistic stylization in a coherent, affectively rich environment. The novelty lies in how these fictionalized archetypes extend traditional CALL agents by embedding language practice within culturally familiar yet emotionally evocative anime-inspired worlds.
3.2. Dialogue System Architecture
The conversational character system in Jouzu was designed to move beyond standard chatbot architectures. Natural conversations between humans and GPT-4 were first leveraged to create QA scripts. Each script was divided into ten sections, with each section containing three to five turns between the user and the LLM. These QA scripts provided the foundation for lessons and informed the training of multimodal text and text-to-speech models. This design highlights a methodological contribution: by grounding conversational scaffolds in pre-authored human–LLM dialogues, Jouzu ensures pedagogical coherence while retaining the flexibility of real-time generation.
3.2.1. Text Generation
During real-time, user-driven interaction, an LLM (e.g., GPT-5) is prompted with contextual cues and instructions to emulate each character’s distinct speaking style. This ensures that user exchanges remain consistent with the intended persona and stylization of the character. The contribution here is in demonstrating how real-time prompting for stylized persona-specific language can extend beyond neutral tutoring into affectively rich, culturally anchored interaction.
3.2.2. Speech Synthesis
Jouzu employs Style-BERT-VITS2 for speech synthesis. Professional Japanese voice actors recorded the QA scripts in each character’s voice, capturing distinct character language and stylistic nuance. These recordings were used to fine-tune models on top of the pretrained Style-BERT-VITS2 JP Extra model. Relevant work shows that Style-BERT-VITS2 JP Extra produces synthesized speech that, on average, shows no significant difference from native Japanese ground truth (rackauckas2025benchmarkingexpressivejapanesecharacter). This integration of professional voice acting with state-of-the-art synthesis contributes to HCI by demonstrating how affectively expressive, near-indistinguishable synthetic speech can be deployed in real-world language learning contexts.
Figure 2. A screenshot from Jouzu featuring the in-text word inspector. The image shows a sent user message with the translation, ”What are you going to do today?” and a response from the character Kitsune. The user has tapped on the sentence, and the flashcard for the Japanese word ”shizen,” translated to ”nature” in English, appears with the furigana, English definition, and romaji spelling of the word.:
3.3. In-text Inspector
Japanese learners, particularly those from non-Chinese language backgrounds, often struggle with reading due to the stark differences from other languages in grammar (boboc2023) and phonology (ohata2004). To address this, Jouzu includes an in-text inspector: during any chat or lesson, users can click on a sentence to trigger flashcards that define each word step-by-step, including readings and pronunciations (see Figure 2). The novelty of this feature lies in its seamless integration of word-level scaffolding into conversational dialogue, offering immediate, interactive support without breaking immersion.
@@ 4. Method
A user study was conducted with 54 participants initially recruited; of these, 52 fully completed all parts of the study, including surveys and app usage. The analysis below focuses on these 52 complete cases to ensure consistency and validity. We examine app usability, learning outcomes, engagement, impact of stylized characters on the learning experience, and comparison of Jouzu to other language learning apps like Duolingo. In addition, we gathered qualitative feedback on user experiences and analyzed interaction patterns between users and conversational agents.
Participants were recruited from a university computer science class. They were then instructed to freely use the app for at least one hour within a period of two weeks. In addition, we asked them to use the conversational agents present in the app. Users completed a survey consisting of multiple choice questions and open responses before (see Appendix A) and after the app (see Appendix B). The initial survey collected demographic and background information from users and presented an optional 10-question multiple choice Japanese quiz with an optional short paragraph response in Japanese (see Appendix C). The post-use survey asked several questions about the user experience, followed by a new 10-question Japanese quiz of the same difficulty level and an open response in Japanese (see Appendix C). Participants were rewarded with 5 assignment points in their university class. Of the 54 recruited participants, 52 fully completed the task.
All participants were students completing this work as part of a classroom assignment. Participation in assessments and activities was voluntary: students who chose not to participate did not incur any penalty, while those who did participate received extra credit. Emails were collected solely for the purpose of linking responses within the study; no identifying details are disclosed in publications, and all data were anonymized prior to analysis. Participation complied with our institution’s policies on classroom-based research activities. In conducting and reporting this work, we adhered to the ACM Policy on Research Involving Human Participants and Subjects, ensuring that participants’ autonomy, privacy, and well-being were respected.
@@ 4.2. Hypotheses and Analysis Design
To examine potential cross-linguistic differences in user experience, participants were grouped by native language: Chinese and non-Chinese speakers. To assess the significance of these cross-linguistic differences, we conducted independent-sample t-tests comparing mean group ratings between Chinese and non-Chinese participants across all six experience categories. We also grouped users by self-reported Japanese proficiency (Complete Beginner, Beginner, Intermediate, Native) to analyze mean ratings by language experience level. A heatmap was generated to visualize rating trends across these categories.
@@ 4.4.1. Multiple Choice
Users were asked ten multiple choice questions before and after use of the app. All questions were optional. The questions are included in Appendix C. Pre- and post-use completion rates and accuracy scores were compared to assess realized learning outcomes in terms of grammar and vocabulary. Ten questions were asked. The first five questions included vocabulary sampled from the characters’ individual language styles. The next five were questions sampled from each Japanese Language Proficiency Test Level (N1 through N5). The pre- and post-tests were different, but were created to be the same difficulty for respondents.
@@ 4.5. Chatbot Interaction Logging
4.6. Usage Statistics
To better understand patterns of engagement, we collected usage data via server-side timestamps. From this data, we derived the following quantitative measures:
  1. (1)
    Average number of chat messages per user: total messages sent per participant, normalized by session count.
  2. (2)
    Average number of sessions per user: discrete usage periods separated by at least 30 minutes of inactivity.
  3. (3)
    Average session duration: computed from the first to last interaction within a session window.
  4. (4)
    Most-used characters: relative frequency of interactions with each stylized agent (Hikari, Kitsune, Samurai).
These measures provided a baseline of overall interaction with the application and allowed us to contextualize survey responses with actual engagement behavior.
4.7. Writing Sample Evaluation (Rubric-Based)
Participants were invited to write a short paragraph in Japanese both before and after using the app. This task was optional and open-ended, allowing users at any level to attempt expressive written output. To evaluate changes in Japanese proficiency, we applied a rubric-based assessment focusing on vocabulary diversity, grammatical complexity, and stylistic features such as character language.
These dimensions follow well-established frameworks in second-language research. Vocabulary diversity has long been recognized as a marker of language development, often measured through indices of lexical richness (malvern2004lexical). Grammatical complexity is similarly central, with studies showing that features like clause chaining and subordination track learner proficiency (lu2010automatic). Finally, we included stylistic and pragmatic appropriateness, drawing on the CEFR’s emphasis on register, politeness, and sociolinguistic control in real communicative settings (council2001cefr). Taken together, these categories reflect the broader complexity–accuracy–fluency (CAF) framework in applied linguistics (housen2012dimensions), ensuring that our rubric aligns with widely accepted standards for assessing learner language.
A near-native Japanese teacher independently scored all pre- and post-use paragraphs on a one, three, or five scale per dimension (higher is better). If no paragraph was written, zero was assigned for each category. D. We deliberately used one/three/five anchors to reduce ambiguity and emphasize categorical distinctions. This contrasts with the 5-point Likert items but improves coding reliability.